OpenStudy (anonymous):

What is the exact value for the expression the square root of 56. − the square root of 14. + the square root of 126.? Simplify if possible. the square root of 14. 4 the square root of 14. 2 the square root of 42. 8 the square root of 42.

OpenStudy (anonymous):

@jigglypuff314 @welshfella @tkhunny

OpenStudy (tkhunny):

Do NOT post without showing your work. \(\sqrt{56} - \sqrt{14} + \sqrt{126}\) Look for perfect squares in the Prime Factorization of each number. Go!

OpenStudy (anonymous):

i don't know how to type out all of the work but i got C

OpenStudy (tkhunny):

Well, then you'll just have to go with that. I think you CAN type it out. 56 = 2*28 = 2*2*14 = 2*2*2*7 Are you SURE you can't do that?